OpenStudy (anonymous):

what is the derivative of (sin2)x?

OpenStudy (anonymous):

not clear what this is \[\sin(2x)\] \[\sin^2(x)\] \[\sin(2) \times x\]?

OpenStudy (anonymous):

sin(2) * x

OpenStudy (anonymous):

well \[\sin(2)\] is a constant, so this is like asking for the derivative of \[.7x\]

OpenStudy (anonymous):

the derivative of \[37x\] is \[.7\] and likewise the derivative of \[\sin(2)\times x\] is \[\sin(2)\]